OpenStudy (anonymous):

How many grams of water can be produced when 65.5 grams of sodium hydroxide reacts with excess sulfuric acid? (4 points) Unbalanced equation: H2SO4 + NaOH → H2O + Na2SO4 @pooja195

pooja195 (pooja195):

H2SO4 + 2 NaOH → 2 H2O + Na2SO4 M(NaOH) = M(Na) + M(O) + M(H) = 23 + 16 + 1 = 40 g/mol M(H2O) = 2M(H) + M(O) = (2 x 1) + 16 = 18 g/mol From the equation : 2 x 40g of NaOH -----> 2 x 18g of H2O 65,5g of NaOH -----> (36 x 65,5) / 80 = 29,5g of H2O

OpenStudy (anonymous):

So how have you come to this conclusion?

OpenStudy (anonymous):

@pooja195 Are you there?

OpenStudy (danjs):